Find the domain of each composite function fog , gof if f(x) =3x+1, g(x)=x2
Answer #1

If f(x) and g(x) has domain as all real number then domain of \large (f\circ g)(x)\;\;\; and \;\;\; (g\circ f)(x)   is all real number . This in interval notation is

\large (-\infty,\infty)

here f(x)=3x+1. And g(x)=x2

these are polynomial function. So domain of both function are all real number .
so domain of fog and gof is

\large \boldsymbol{domain=(-\infty,\infty)}